OpenStudy (anonymous):

The arithmetic mean of three numbers is 147. If the mean of the first two numbers is 141, what is the third number?

OpenStudy (anonymous):

(x+y+z)/3=147 (x+y+z)=441 (x+y)/2=141 (x+y)=282 z=441-(x+y) z=441-282=159